3. You are given the following information about the economy:

autonomous consumption = $300 billion       planned investment = $300 billion

government spending = $500 billion             mpc = .8

imports = $200 billion                                    exports = $500 billion                       

a. Using the values above, what is the equation for the consumption function?

b. Using the values above, what is the income/spending multiplier?

c. What is the value of Net Exports?

d. Is there a trade surplus or deficit? Of how much?

e. Using these values, what is the equation of aggregate demand?

f. What is the equilibrium level of Y?

g. If full employment Y = $6000 billion, is there a recessionary or inflationary gap?

h. If full employment Y = $6000 billion, what is the GDP gap?

i. What kind of fiscal policy (expansionary or contractionary) could help the economy close the gap?

j. If you decided to change government spending, how much would you change G to close the GDP gap?

k. If the value of the dollar depreciated, how would this affect net exports? AD?
Answer #1

a) C = Autonomous consumption + MPC x Disposable income = 300 billion + 0.8(Y - T)

b) Multiplier = 1/(1 - MPC) = 1/(1 - 0.8) = 1/0.2 = 5

c) Net exports = Export - Import = 500 billion - 200 billion = 300 billion

d) Trade surplus = Export - Import = 500 billion - 200 billion = 300 billion

e) AD = C + I + G + Exports - Imports

AD = 300 billion + 0.8(Y - T) + 300 billion + 500 billion + 500 billion - 200 billion

AD = 1400 billion + 0.8(Y - 0) = 1400 billion + 0.8Y

f) At equilibrium; Y = AD

Y = 1400 billion + 0.8Y

Y - 0.8Y = 1400 billion

0.2Y = 1400 billion

Y = 7000 billion